What is the size of an angle whose complement is 25%of its supplement
Answer by mananth(16946)

let angle be x
supplement = 180-x
complement = 90-x
...
25% = 1/4
..
90-x = 1/4 * (180-x)
360-4x = 180-x
4x-x=360-180
3x=180
x= 60 deg
...
m.ananth@hotmail.ca